OpenStudy (anonymous):

Find the equation of the line in slope-intercept form containing the points (6, -1) and (-3, 2).

pooja195 (pooja195):

We need to find the slope first

pooja195 (pooja195):

\[\huge~\frac{ y_2-y_1 }{ x_2-x_1 }\]

pooja195 (pooja195):

(x1,y1) (x2,y2)

OpenStudy (anonymous):

3/-3

OpenStudy (anonymous):

i think?

OpenStudy (mathmate):

(x1,y1)=(-3,2) (x2,y2)=(6,-1) So (y2-y1)/(x2-x1) = (-1-2)/(6-(-3)) = ?

OpenStudy (mathmate):

It is always a good idea to write down the two points one on top of the other, as @pooja195 suggested earlier. This way, it's easier to figure out the numbers without making a mistake.

OpenStudy (anonymous):

3/-3

OpenStudy (anonymous):

wait 1/-3

OpenStudy (mathmate):

Excellent, so that's your slope, 1/-3 is the same as -1/3. Now use the point slope form to get the equation. You can use P1=(x1,y1)=(-3,2) as your point, and slope is a=-1/3. The equation is then (y-y1) = a(x-x1)

OpenStudy (anonymous):

y= 1/3 x-1?

OpenStudy (mathmate):

You are close, but got signs wrong somewhere.

OpenStudy (anonymous):

y=-1/3 x+1?

OpenStudy (mathmate):

That's it, excellent! Is everything clear?

OpenStudy (anonymous):

yes thank you so much!!

OpenStudy (mathmate):

You're welcome! @pooja195 started you off with the right foot!